OpenStudy (anonymous):

can x^3+7x^2+8x-16 be simplified any further?

jimthompson5910 (jim_thompson5910):

no because there are no like terms if you had x^3 and 7x^3, you can combine them to get x^3 + 7x^3 = 8x^3

OpenStudy (anonymous):

can you do the thing (idk what its called) where you find a number that when added equals 8 but when multiplied equals -16?

jimthompson5910 (jim_thompson5910):

that only works with quadratics

jimthompson5910 (jim_thompson5910):

and you're thinking of factoring

OpenStudy (anonymous):

ooh yeah that's right

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

do long division

OpenStudy (anonymous):

where would I do that

OpenStudy (anonymous):

@jim_thompson5910 can you check my answer? I had to multiply (x^2+3x-4)(x+4) and the final answer I got was x^3+7x^2+8x-16

jimthompson5910 (jim_thompson5910):

let me check

jimthompson5910 (jim_thompson5910):

I'm getting the same thing.

OpenStudy (anonymous):

I thought it was right I was just making sure thanks again

OpenStudy (anonymous):

i did long division (x-1)(x^2+8x+16)=x^3+7x^2+8x+16

OpenStudy (anonymous):

now you could simplify it more x^2+8x+16=(x+4)(x+4) the answer is (x+4)(x+4)(x-1)

OpenStudy (anonymous):

i hope it is clear now